OpenStudy (anonymous):

Find the derivative of tan(x)/x^3

OpenStudy (anonymous):

quotient rule

OpenStudy (shamil98):

quotient rule

OpenStudy (shamil98):

\[\huge \frac{ f'g - g'f }{ g^2}\]

OpenStudy (anonymous):

We use u and v in Australian education Interesting

OpenStudy (shamil98):

it's the same result, but then i guess it would be: \[\huge \frac{ u'v-v'u }{ v^2 }\]

OpenStudy (anonymous):

We do this and then it's easier to remember where the letters go becuase they make a v in the fraction. It is indeed the same result \[\frac{ vu'-uv' }{ v^2 }\]